OpenStudy (anonymous):

find the instantaneous rate of change of the function y=3x-5cos^3(pix)

OpenStudy (anonymous):

at a given number or just find the derivative?

OpenStudy (anonymous):

just the derivative

OpenStudy (anonymous):

chain rule for this one

OpenStudy (anonymous):

\[y'=5-15\cos^2(\pi x)\times (-\sin(\pi x))\times \pi\]which you can clean up a bit

OpenStudy (anonymous):

typo there, that first 5 should be a 3

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw